OpenStudy (ninja):

What is the solution to the equation log2 4 – log2 x = 3 ?

OpenStudy (lgbasallote):

log2(4/x) = 3 2^3 =4/x 8 = 4/x 8x = 4 x = 4/8 = 1/2

OpenStudy (anonymous):

1/2